OpenStudy (anonymous):

Algebra II. Equations with Paretheses. A) 12(x -6) = 2(x - 9) B) 4(7 - 3x) = 7(4 - 2x) C) 7(x - 2) + 3(x + 2) = 5(x - 6) D) 4(2x -3) + 8(x - 4) = 2(2x + 6) E) 4(7x - 1) = 3(2x + 1) - 4(x + 6) F) 12(2x - 8) = 4(5 - 2x) + 3(7 - 3x) G) 3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1) Any help you could provide to giving or helping me find these answers would be lovely! :)

OpenStudy (anonymous):

12x-72=2x-18 10x=90 x=9 Just multipy each one out and then add or subtract like terms to get x on one side and constants on the other then solve for x

OpenStudy (anonymous):

This problem look more like an Algebra I problem. I will do (E)..then you work on the rest. All same type. 4(7x-1) = 3(2x+1) - 4(x+6) Use Distributive Law, to get rid of parenthesis, 28x - 4 = 6x + 3 - 4x - 24 28x - 4 = 2x - 21 26x =-17 x = -17/26

OpenStudy (anonymous):

28x - 4 = 6x + 3 - 4x - 24 is x=17/26